A simple pendulum is suspended from the ceiling of a car accelerating uniformly on a horizontal road.If the acceleration is a and length of the pendulum is l,find the time period of small oscillation about the mean position.

Grade:12

3 Answers

Chetan Mandayam Nayakar
11 years ago

T=2pi√(l/(√(a2+g2)))

Yogita Bang
11 years ago

The car is accelerating in horizontal direction with acceleration a. The accleration due to gravity is g in downward direction.

Resultant acceleration a'' = (a2 + g2)1/2

Time period T=2pi(L/a'')1/2 =2pi{L/(a2+g2)1/2}

 

 

Kushagra Madhukar
3 years ago
Dear student,
Please find the solution to your problem.
 
The car is accelerating in horizontal direction with acceleration a. The accleration due to gravity is g in downward direction.
Resultant acceleration a'' = (a2 + g2)
Time period T = 2π(L/a'') = 2π{L/(a2+g2)}
 
Thanks and regards,
Kushagra